OpenStudy (anonymous):

A customer went to a garden shop and bought some potting soil for 12.50 and 5 shrubs.The total bill was 62.50. Write and solve and equation to find the price of each shrub

OpenStudy (anonymous):

s = # of shrubs 12.50 + 5s = 62.50 -- subtract 12.50 from both sides 5s = 62.50 - 12.50 5s = 50.00 -- now divide both sides by 5 s = 10.00 check.. 12.50 + 5(10) = 62.50 12.50 + 50.00 = 62.50 62.50 = 62.50 (correct) answer is : $10 per shrub